The x and y components of a certain force are measured and found to be Fx = (68 ± 3)N and Fy =(42 ± 2)N, respectively. Calculate the magnitude of this force? expressing your result to the proper number of significant figures.
Answer #1

Ex = 680 fy= 42N. magnitude , f= JFx²4 Fy2 = 568²442² - 79.925 N manimum value of fx = 68+3 = Fy = 42+2 = 70N UUN Maremum val